Shift report 20011031 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Oct 31 2001 - 00:10:46 EST

  • Next message: Apache: "Shift report 20011031 00:00-08:00"

    ------------------------------------------------------------
    Date: Wed Oct 31 00:10:46 2001, Shift: 16:00-24:00
    Supervisor: rw, with: 
    Log book#10 pages 75 - 76
    
    Shift Summary:
    ------------------------------------------------------------
    1. At 21:45 a good storage was made. blue/yellow=29/27*10**9
    ZDC 900-1000. (Star and Phenix are 1400-1500) An additional steering request did not 
    improve the situation more than 5%.
    2.
    	run 5589   TPC Pedestals
                5590    DVM run
      	    5591   Pedestals
    	    5592    Data (FS @ 8deg. 1/5A MRS @45deg 4B
    		   Trigger 6 200K.
     	    5593    Same as the previous run. running
    3. T5 2nd plane HV off. T3 currents 53, 47,41 micro Amp.
    No other problems are noticed.   
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Oct 31 2001 - 00:11:12 EST